solve it root x + y= 11 


and x + root y =7 

Grade:12

7 Answers

Aman Bansal
12 years ago

Dear Keshav,

Root X + y = 7
Root Y = 11

Thus, x + y = 7^2 (bring square root over, becomes squared)
and y = 11^2 (again square root over)

So we then know that y = 121
Sub it in:

x + y = 49
x +121=49
x =49 - 121
x = -72

So we now know y = 121 and x = -72

Best Of luck

Plz Approve the answer...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….
So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

keshav kr
10 years ago
wrong ans by iitian
Shrey
10 years ago
Equation 1- Root(x+y)=11 Equation 2- x+root y=7 squaring both sides of equation 1 we get (x+y)=121->1. (x+root y)=7->2. Subtracting 1. From 2. Y-root y=114 root y(root y-1)=114 so either root y=114 or root y =114+1=115 so root y=114 or 115 squaring both sides y=(114)^2 or(115)^2 similarly find x by putting value of y in any of the equations. Please approve my answer by clicking approved if you liked it.
Ranvir Sharma
8 years ago
x=4
y=9
Vinod yadav
6 years ago
Fist equation was just root x + y(separate) = 11.. so after squaring,how can u have just x + y on LHS...it is not the whole root of x + y.... Such a silly mistake...didn`t expect it on this forum...
Meet
6 years ago
Firstly take √x=m and take √y=t and the modified equation is m+t^2=11 and m^2+t=9 after solving these two equation u will get x=4 and y=9.
Vinod yadav
6 years ago
Just writing that after solving you will get the answer as 4& 9 is not sufficient. Show the steps and justify ....